OpenStudy (anonymous):

anti derivative of (x+1)/ (sqrt (1-x^2))

OpenStudy (irishboy123):

note that \(\frac{d}{dx} \sqrt{1-x^2} = \frac{1}{2} (1-x^2)^{-\frac{1}{2}}(-2x) = ???\) with a bit of fiddling, that deals with first part of the integrand's numerator for the other part of the numerator, you should try *the* trig sub

OpenStudy (anonymous):

can you walk me through that or just do it so i can see it

OpenStudy (irishboy123):

nah, there's only so much i can do before I get another email saying i did someone's work for them.....:-( but we/you can tidy this up..... \[\frac{d}{dx} \sqrt{1-x^2} = \frac{1}{2} (1-x^2)^{-\frac{1}{2}}(-2x) = \dfrac{-x}{\sqrt{1-x^2}}\] and you are integrating \(\dfrac{\color{red}{x}+1}{\sqrt {1-x^2}}\) you really should now see what you are left with. good luck!! if no-one helps you, tag me....